OpenStudy (anonymous):

Find the average rate of change of f(x)=2x^3-2x^2-3 from x=1 to x=3. Not sure what to do with the 2x^3 and the 2x^2 -thats where i get lost...

OpenStudy (cruffo):

average rate of change = slope of secant line = change in y / change in x. You need to find f(1) = "y1" and f(3) = "y2". Then find the slope of the line connecting the two points (1, y1) and (3,y2).

OpenStudy (anonymous):

\[f(3)=2\times 3^3-2\times 3^2-3=33\] \[f(1)=2-2-3=-3\]

OpenStudy (anonymous):

thats as far as i have gotten

myininaya (myininaya):

\[\frac{33-(-3)}{3-1}=\frac{36}{2}=\frac{18}{1}=18\]

OpenStudy (anonymous):

\[\frac{33-(-3)}{3-1}=\frac{36}{2}=18\]

OpenStudy (anonymous):

lol

myininaya (myininaya):

i win satellite

OpenStudy (anonymous):

fastest latex in the west

myininaya (myininaya):

i even put 18/1 in there (not needed though lol)

OpenStudy (anonymous):

omg lol